PFI-ZEKE SPECTROSCOPY OF Be_{2}O

\underline{I. O. ANTONOV}, M. C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The PFI-ZEKE spectrum of Be_{2}O^{+} ion was recorded to obtain accurate ionization energy for Be_{2}O and vibrational frequency data for Be_{2}O^{+}. A short progression of the bending mode was observed, which included bands involving odd changes in the vibrational quanta. This observation indicates that the ion and/or the intermediate state used in the excitation process deviate from D_{∞}h symmetry.